OpenStudy (anonymous):

why is the line y=-x an asymptote for the graph of y=(1-x^2)/x

OpenStudy (anonymous):

Because the line in the graph draws towards y=-x but never touches it.

jimthompson5910 (jim_thompson5910):

Divide (1-x^2) by x to get (1-x^2)/x = 1/x - (x^2)/x = 1/x - x So we get 1/x - x Now if x approaches infinity, then 1/x will get very very small (and will approach zero) So when x approaches infinity, 1/x will basically go away. This leaves us with -x So the slant (or oblique) asymptote is y = -x

OpenStudy (anonymous):

thank you!